June 1 balance 28,000

Materials..................................................................................... 120,000

Direct Labor......................................................................................79,500

Overhead...........................................................................................97,000

Raw materials...……………………………………………………………….$45,000

Work in process...………………………………………………………… 20,000

Finished goods........................................................................................40,000

June 30 balance............??????

The company uses a plant wide rate to apply overhead. The overhead rate was based on an estimated over head cost of $640,000 and an estimated level of activity of 32,000 direct labor hours. The June work in progress inventory consisted of 5,000 pounds with 16,000 in materials cost and 12,000 in conversion cost. The June 1 work in progress inventory was 100% complete with respect to materials and 50% complete with respect to conversion. During June, 37,500 pounds were started into production. The June 30 work in progress inventory consisted of 8,000 pounds that were 100% complete with respect to materials and 40% complete with respect to conversion.

1. Prepare the journal entry to record the overhead cost applied to production.

2. Compute the equivalent units of production for materials.

3. Compute the equivalent units of production for conversion.

4. What is the amount of of the cost of beginning work in progress inventory plus the cost added during the period for conversion?

5. What is the cost per equivalent unit for materials?

6. What is the cost per equivalent unit for conversion?

7. What is the cost of ending work in progress for materials?

8. What is the cost of materials transferred to finish goods?

9. What is the amount of conversion cost transferred to finished goods?

10. Prepare the journal entry to record the transfer of costs from Work in progress to Finished goods.

11. What is the total cost to be accounted for? what is the total cost accounted for?
Answer #1

1.

Account Titles Debit Credit
Work in Process Inventory $        97,000
       Manufacturing Overhead $        97,000

2. Equivalent units of Production = Units transferred + Ending WIP x Completion
Units transferred = 5000+37500-8000 = 34500
Equivalent units
Material = 34500 + 8000 x 100% = 42500 pounds

3.
Conversion = 34500 + 8000 x 40% = 37700 pounds

4.

Costs to be accounted for
Cost of beginning work in process $        28,000
Cost added during production $     296,500
Total Costs to be accounted for $     324,500

5. Cost per equivalent unit = (Beginning Cost + Cost added) / Equivalent units of production
Material = ($16000+12000)/42500 = $3.20 per pound

6.
Conversion = ($12000+79500+97000)/37700 = $5 per pound

7. Cost of ending work in process for materials = 8000 x 100% x $3.20 = $25600

8. Cost of materials transferred to finished goods = 34500 x $3.20 = $110400

9. Cost of conversion transferred to finished goods = 34500 x $5 = $172500

10.

Account Titles Debit Credit
Finished Goods Inventory $     282,900
       Work in Process Inventory $     282,900

11.

Costs to be accounted for
Cost of beginning work in process $        28,000
Cost added during production $     296,500
Total Costs to be accounted for $     324,500
Costs accounted for as follows
Cost of Ending work in process $        41,600
Cost of Units transferred $     282,900
Total Costs accounted for $     324,500